Write the explicit formula for the arithmetic sequence where t underscore 7 =121 and t underscore 15 =193 .

Answer by stanbon(75887)

Write the explicit formula for the arithmetic sequence where t underscore 7 =121 and t underscore 15 =193 .
------
t(7) = a + 6d = 121
t(15) = a + 14d = 193
-----------------------------
Subtract and solve for "d"::
8d = 72
d = 9
-----
Solve for "a"::
a + 6*9 = 121
a = 121- 54
a = 67
------------
Ans: t(n) = 67 + (n-1)*9 = 67 - 9 + 9n = 58 + 9n
-------------
Cheers,
Stan H.